OpenStudy (anonymous):

Solve the polynomial: 0= x2 + 2x +5. Be sure to put down your work in your answer and EXPLAIN HOW YOU DID IT!!!! @Hotchellerae21

OpenStudy (anonymous):

i used the quadratic formula

OpenStudy (anonymous):

What is the quadratic formula @satellite73?

OpenStudy (anonymous):

or better yet, complete the square it will be easier

OpenStudy (anonymous):

Thank you so much but please tell me what the quadratic formula is!!!!!! Thank you so much.

OpenStudy (nerdgirl):

will one of you guys please come help me in my algebra 2 problem when you're done here?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Yeah sure @Nerdgirl

OpenStudy (anonymous):

\[x^2+2x+5=0\] the zeros will be complex subtract 5 get \[x^2+2x=-5\] complete the square get \[(x+1)^2=-5+1\\ (x- +1)^2=-4\] take the square root \[x+1=\pm\sqrt{-4}\\ x+1=2i\]

OpenStudy (anonymous):

subtract 1 and \[x=-1\pm2i\]